OpenStudy (anonymous):

3(3a+3)+6=81 (Show Work)

OpenStudy (anonymous):

3(3a+3)+6=81 9a+9+6=81 9a+15=81 9a=81-15 9a=66 a=66/9

OpenStudy (anonymous):

3(3a+3)+6=81 9a+9+6=81 9a+15=81 -15 -15 9a=66 a=66/9